OpenStudy (anonymous):

Find the least common denominator of 7/6p and 15/4p-8

OpenStudy (anonymous):

\[\frac{ 7 }{ 6p }and \frac{ 15 }{ 4p-8 }\]

OpenStudy (anonymous):

find lcd

OpenStudy (mathstudent55):

6p = 2 x 3 x p 4p - 8 = 2(2p - 4) Of all the factors from both denominators, you need one of each: 2, 3, p, 2p -4

OpenStudy (anonymous):

ok got it..whats the next step?

OpenStudy (mathstudent55):

multiply all factors together and that's your LCD

OpenStudy (mathstudent55):

2(3)p(2p - 4) = LCD

OpenStudy (anonymous):

the answer is 12p(p-2) just dont know how they got it

OpenStudy (mathstudent55):

Sorry, 2p - 4 can be factored to 2(p - 2)

OpenStudy (mathstudent55):

So factors of first denom are 2,3,p Factors of second denom are 2,2,p - 2 Since you have one 2 in first and 2 2's in second, you need bioth 2's. You have one 3 in first take it too. then you need the p in first and the p - 2 in second. So altogether you need 2 x 2 x 3 x (p - 2)

OpenStudy (mathstudent55):

Last line should read So altogether you need 2 x 2 x 3 x p x (p - 2)

OpenStudy (mathstudent55):

When you multiply that together you get 12p(p - 2)

OpenStudy (anonymous):

ok thanks

OpenStudy (mathstudent55):

wlcm